# In the Network of the Conclave: Social Network Analysis and the Making of a Pope

### v1

@@ 6 Discussion
Needless to say, this approach is deeply rooted in a long tradition of studies and research. From Katz,’s (1957) diffusion thresholds to Freeman,’s (1977) betweenness to Granovetter, (1973) weak ties, the literature on networks has long showed that position matters. The conclave reminds us that classical tools remain potent when applied with care. By projecting a bipartite affiliation structure and layering a sacramental lineage onto it, we have shown how simple techniques illuminate a domain that, at first glance, seems empirically intractable. While our study sheds light on the structural underpinning associated with being papabile within the College of Cardinals, several limitations deserve careful attention. First, our network reconstruction is based entirely on formal ties—co-membership in ecclesiastical bodies and consecration lineages. These ties offer a meaningful proxy of important relationships, but they inevitably miss the informal social ties among cardinals. Friendship ties, informal mentoring relationships, and personal rivalries or negative ties (Kilduff et al.,, 2024; Labianca and Brass,, 2006), which could play a key role in decision making, remain outside the scope of our investigation. Second, we do not capture perceived cognitive social structures, meaning that we don’t know how cardinals perceive the network they are embedded in. Research has shown that actors often rely on subjective network maps when making strategic decisions (Krackhardt,, 1987), and these maps can diverge sharply from objective structure. A cardinal may appear central in our reconstruction, yet be perceived as peripheral by his peers, or vice versa (Iorio,, 2022). Without access to first-hand perceptions, interviews, or internal documents, we cannot account for this cognitive dimension of influence. Third, our approach is cross-sectional and static. It does not model how the network evolves in the lead-up to the conclave or preferences shift across successive ballots. Nor do we account for external interventions (e.g., regional lobbying) that may reconfigure priorities in real time. Finally, our analysis is bounded by observability and survivorship: we analyze the College as it existed in April 2025, but the relational dynamics that matter most may have formed over decades of prior interaction, some of which leave no public trace. Future research may explore ways to approximate these hidden histories—through qualitative fieldwork, ecclesiastical archives, or digital records of public appearances.
In conclusion, we would like to reiterate one more time that our analysis stops deliberately short of prediction. What we can offer is a structural map that narrows the plausible space of outcomes and could help explain why certain candidacies crystallize while others fade. Methodologically, the work is constrained by reliance on public records: informal advice ties, spiritual friendships, and personal antipathies remain outside our map. Future research could triangulate archival sources, ethnographic interviews, or digital traces (e.g., homily co‐participation) to approximate these softer layers. Dynamic network models—estimating how support percolates ballot by ballot—are an obvious next frontier, though they will require creative approaches to data collection.
@@ 7 Conclusion
The broader takeaway of our study is straightforward: no serious analysis of future conclaves can afford to ignore network structure. Whether scholars pursue qualitative thick description, statistical modelling, or computational simulation, the relational scaffolding of the College must serve as the baseline upon which other factors—doctrinal, regional, theological—operate. The conclave is not merely a gathering of isolated cardinal electors; it is an interaction network with a long memory and a short time horizon. To attempt to decipher its logic we must put a relational perspective back at the center of the stage.
@@ Appendix A Appendix: Online Article
As mentioned, this data collection was based only on apical figures of the Roman Curia and included a historical view of comembership in collegial bodies, assuming ties persist over time. As we explain, this analysis also incorporats a small, selected number of purely informal ties based on interpersonal bonds and relationships, as identified through journalistic reports that we accessed. These informal ties play a minimal role in the College of Cardinals network given their limited number.

### v2

@@ 6 Discussion
While our study sheds light on the structural underpinning associated with being papabile within the College of Cardinals, several limitations deserve careful attention. First, our network reconstruction is based entirely on formal ties—co-membership in ecclesiastical bodies and consecration lineages. These ties offer a meaningful proxy of important relationships, but they inevitably miss the informal social ties among cardinals. Friendship ties, informal mentoring relationships, and personal rivalries or negative ties (Kilduff et al.,, 2024; Labianca and Brass,, 2006), which could play a key role in decision making, remain outside the scope of our investigation. Second, our arguments linking the prominence of Cardinal Prevost to his position in the Vatican network assume that co-membership and co-consecration ties are visible to the College of Cardinals, while opaque for external observers. Specifically, within the College of Cardinals, there is high visibility of the roles underlying co-membership ties and a clear awareness of consecration links. However, while visible, interpersonal ties may or may be not perceived accurately by cardinals. In this regard, we do not capture perceived cognitive social structures, meaning that we don’t know how cardinals perceive the network they are embedded in. Research has shown that actors often rely on subjective network maps when making strategic decisions (Krackhardt,, 1987), and these maps can diverge sharply from objective structure. A cardinal may appear central in our reconstruction, yet be perceived as peripheral by his peers, or vice versa (Iorio,, 2022). Without access to first-hand perceptions, interviews, or internal documents, we cannot account for this cognitive dimension of influence (Aslarus et al.,, 2025). Third, our approach is cross-sectional and static. It does not model how the network evolves in the lead-up to the conclave or preferences shift across successive ballots. Nor do we account for external interventions (e.g., regional lobbying) that may reconfigure priorities in real time. Finally, our analysis is bounded by observability and survivorship: we analyze the College as it existed in April 2025, but the relational dynamics that matter most may have formed over decades of prior interaction, some of which leave no public trace. Future research may explore ways to approximate these hidden histories—through qualitative fieldwork, ecclesiastical archives, or digital records of public appearances.
Finally, while our analysis deliberately stops short of making predictions, it offers a structural map that narrows the range of plausible outcomes—and helps illuminate why some candidacies crystallize while others fade.
@@ 7 Conclusion
The broader takeaway of our study is straightforward: no serious analysis of future conclaves can afford to ignore network structure. Whether scholars pursue qualitative thick description, statistical modelling, or computational simulation, the relational scaffolding of the College must serve as the baseline upon which other factors—doctrinal, regional, and theological—operate. The conclave is not merely a gathering of isolated cardinal electors; it is an interaction network with a long memory and a short time horizon. To attempt to decipher its logic, we must put a relational perspective back at the center of the stage.
@@ Appendix A Appendix: Online Article
As mentioned, this data collection was based only on top figures of the Roman Curia and included a historical view of comembership in collegial bodies, assuming ties persist over time. As we explain, this analysis also incorporats a small, selected number of purely informal ties based on interpersonal bonds and relationships, as identified through journalistic reports that we accessed. These informal ties play a minimal role in the College of Cardinals network given their limited number.